# Compositional disorder in a multicomponent non-reciprocal mixture: stability and patterns

### v2

@@ III Spinodal
We run $50$ realizations of our simulation for $N=100$ components to obtain the results in Fig. 2.
Figure 3: Boundary enclosing the eigenvalues in the complex plane (a) The boundary and the spectrum of a large random matrix with diagonal disorder drawn from the bimodal distribution, see Eq. (13) for $d_{2}=-d_{1}=d$. (b) The diagonal elements of $\mathbb{C}$ are drawn from a uniform distribution where $d_{-}=1$ and $d_{+}=2$. (c) The diagonal disorder is drawn from an exponential distribution with $d=0.6$. (d) The spectrum and the boundary of support for a large random matrix with diagonal elements from a beta distribution with $a=2$ and $b=5$. For all examples, the random matrix is of the size $\text{N}=1000$ and $\chi=1,\alpha=0.5$. (e) The diagonal disorder are drawn from a trimodal distribution with $d_{1}=0.5$, $d_{2}=2$, $d_{3}=5$, $p_{1}=\frac{1}{2}$ and $p_{2}=\frac{1}{5}$ the random matrix is of a size $N=2000$ and $\chi=1$.
Figure 4: The slope $m$ and the intercept $l_{0}$ for different types of compositional disorder. The slope $m$, referred to in Eq. (8) and (IV) for the definition, is calculated for the distributions indicated in the figure. The heatmap of $m$ for the two parameters in (a) and (d) and its variation with the width $d$ in (b) and (c) shows that it is always less than zero. Both the uniform and exponential distributions approach the delta function for $d\to 0$ so that $m$ approaches $-1$ in panels (b) and (c). Panels (e) and (f) show that $l_{0}$ can be positive or negative, meaning that $\Theta_{\rm sp}$ for the passive system can be higher or lower for the corresponding system with identical mean densities.
IV Eigenvalue spectrum
In this section, we will explore four different distributions which are either bounded or unbounded, and which assume either discrete or continuous values to further explore the results in the previous section. The boundary enclosing the spectrum of complex eigenvalues is well approximated by a continuous curve when $N\to\infty$, see Fig. 3 for some examples. In Fig. 3, the markers denote eigenvalues of a single realization of the matrix $\mathbb{D}$ while the boundary has been calculated using methods described in [66], which we will briefly describe here. The spectral distribution $\rho$ of the eigenvalues $\lambda_{j}$ of $\mathbb{D}$ can be written formally as
@@ IV Eigenvalue spectrum
The boundary of the spectrum where $\rho(z)$ is be called $S$ and is given by the points $z\equiv(x,y)$ in the complex plane. The starting point of our analysis is the coupled equations determining the contour of points $(x,y)$ in the complex plane that envelopes the spectrum of $\mathbb{D}$. To do so, it is essential to introduce another quantity $\mbox{Re}(g)$ the physical meaning of which has been mentioned in the Appendix VII.
$\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{\bar{\phi}-x+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}}{\left((\bar{\phi}-x)+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}\right)^{2}+{y^{2}}/{(2\alpha)^{2}}}$
$=-\operatorname{Re}\left(g\right),$ (11)
and
$\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{\sigma^{2}}{\left((\bar{\phi}-x)+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}\right)^{2}+{y^{2}}/{(2\alpha)^{2}}}$
$=1.$ (12)
For $P(\bar{\phi})=\delta(\bar{\phi}-\phi_{0})$, the eigenvalues are distributed uniformly with an ellipse with major and minor axes equal to $1-\alpha^{2}$ and $\alpha^{2}$ respectively, such that $\mbox{Re}[\lambda_{0}]=1-\alpha^{2}$ as can be verified readily from Eqs. (11) and (12).
@@
As seen in Fig. 3 (a), the bimodal effectively splits the system into two groups containing $pN$ and $(1-p)N$ species each. For a fixed $p$ and $d$ close to 00 we find that the boundary of the spectrum simply elongates in the direction of the real axis. As $d$ is increased further, a symmetric fissure appears in the boundary before it is symmetrically split into two contours. The numerical results match well with the contour determined from solving Eqs. (11) and (12).
The exponential distribution is unbounded and associated with a single parameter $d$ that determines the upper-limit of the values that are sampled with comparatively higher probability.
@@
The boundary in Fig. 3 (b) reflects the asymmetry of the distribution and shows a tail that reflects the large deviations due to a few species with high mean composition.
@@
where $N_{a,b}$ is the normalization constant and $B$ is the Beta function.
We next motivate the general relation in Eq. (8) by a closer examination of the structure of the coupled equations Eqs. (11) and (12) and verify it by numerical calculations using the distributions introduced above. To do so, we define new variables
@@
using which we can re-write Eq. (11) as
$\mbox{Re}(g)=-\int\mbox{d}\bar{\phi}\,P(\bar{\phi})\frac{(\bar{\phi}-\tilde{x})}{(\bar{\phi}-\tilde{x})^{2}+\tilde{y}^{2}}\equiv f_{\tilde{y}}(\tilde{x}),$ (18)
where the equivalence defines the function $f_{\tilde{y}}$ whose form depends on $\tilde{y}$. Similarly, Eq. (12) can be re-expressed using a function $g_{\tilde{y}}$
$\frac{1}{\sigma^{2}}=\int\mbox{d}\bar{\phi}\,P(\bar{\phi})\frac{1}{(\bar{\phi}-\tilde{x})^{2}+\tilde{y}^{2}}\equiv g_{\tilde{y}}(\tilde{x}),$ (19)
where the subscript $y$ denotes the value of $\tilde{y}$ at which the integral Eq. (19) is evaluated. Eq. (19) can be inverted to obtain
$\tilde{x}$ $=$ $g_{\tilde{y}}\left(-\sigma^{-2}\right),$ (20)
As seen in Fig. 3 (a), $y$ Similarly inverting Eq. (18), we have $\mbox{Re}(g)=f^{-1}_{\tilde{y}}(\tilde{x})$, which can be used together with Eq. (20) to obtain
$\mbox{Re}(g)$ $=$ $f^{-1}_{\tilde{y}}\left[g_{\tilde{y}}\left(-\sigma^{-2}\right)\right].$ (21)
for a given $\tilde{y}=y/\alpha^{2}$, we can in principle determine the functions $f_{\tilde{y}}$ and $g_{\tilde{y}}$ which can be used to determine $\tilde{x}$ and $\mbox{Re}(g)$ to obtain $x$ as
$x=g_{\tilde{y}}\left(-\frac{1}{\sigma^{2}}\right)+\mbox{Re}(g)(1-2\alpha^{2})\sigma^{2}.$ (22)
From Eq. (17) we deduce that the curves $S$ for two different values of $\alpha$, the $y$ values are proportional. For $\tilde{y}=0$, i.e., the points at which $S$ intersects the $x$ axis are the same for $S$ for any value of $\alpha$. This means that the topology of $S$ is determined by $\{D\}$ only and does not depend on $\alpha$ at all, as illustrated using the bimodal distribution in panel (e) of Fig. 3.
By the definition of $\lambda_{0}$, we have using Eq. (22)
$\lambda_{0}=L_{0}-m\alpha^{2}.$ (23)
where
$L_{0}=g_{0}\left(-\sigma^{-2}\right)+f^{-1}_{0}\left[g_{0}\left(-\sigma^{-2}\right)\right],$
$m=2\sigma^{2}\mbox{Re}(g)=2\sigma^{2}f^{-1}_{0}\left[g_{0}\left(-\sigma^{-2}\right)\right].$ (24)
We note that $\alpha=1/\sqrt{2}$ is a special point because at this point $\tilde{x}=x$, and it is particularly easy to carry out the algebraic operations described above, as shown with a few examples in the Appendix VII. It is also relatively easy to discuss the sign of $m$ at this value of $\alpha$. The diagonal disorders that we consider are always positive quantities and thus bounded by zero from below. From Eq. (18) it is clear that at the extreme value $\lambda_{0}$ of $\tilde{x}=x$, the sign of the integrand is determined by $(x-\bar{\phi})$, which is always negative in the domain of $\bar{\phi}$.
To test that the predictions are correct, we plot the curvature $m$ by varying the parameters $\{D\}$ for the distribution shown in Fig. 1. For each value indicated in the axes of the panels in Fig. 4, we calculate $\lambda_{0}$ at two values of $\alpha^{2}$ and calculate the curvature and verify that it is indeed positive everywhere.
$m=\frac{\lambda_{01}-\lambda_{02}}{\alpha^{2}_{2}-\alpha^{2}_{1}}.$ (25)
Considering again the point $\alpha=1/\sqrt{2}$, we can argue that the topology of the curve is restrained in a particular way.
$\frac{x(\alpha_{1})-x(\alpha_{2})}{\alpha_{2}^{2}-\alpha_{1}^{2}}=2\mbox{Re}(g)(\tilde{y})\sigma^{2}$ (26)
@@
Eq. (22) suggests that the curve $S$ can be calculated at all values of $\alpha$ if we calculate it at $\alpha=1/\sqrt{2}$. For the bimodal distribution, the curve in $S$ splits into two above a threshold value of $d$, meaning that $\tilde{y}$ vanishes at four points on the $x-$axis. The topology of the boundary is preserved for all values of $\alpha$ as shown in Fig. 3(e) for a trimodal distribution where the $\{D\}$ are fixed and $\alpha$ only is varied. The spectrum for the values indicated at the values of alpha indeed follow the insight from theory.
The points at which $S$ intersects the real axis are unchanged for any value of $\alpha$.
@@
Figure 5: Statistics of $\lambda_{0}.$ In the limiting case of infinite conserved densities, $\lambda_{0}$ is real. For any finite sized system, $\langle\lambda_{0}\rangle$ averaged over ensembles of $\mathbb{D}$, is in fact complex. (a) The imaginary and reals parts, $\mbox{Im}(\lambda_{0})$ and $\mbox{Re}(\lambda_{0})$ are plotted as a function of $\alpha$, encoded in the colour of the marker. The error bar indicates the variance. At each value of $\alpha$, we consider $10^{4}$ realizations of $\mathbb{D}$ for $N=300$. From the plots, it is clear that the variability in $\mbox{Im}(\lambda_{0})$ depends on the type of compositional disorder. Notice also that the variations are increased compared to the uniform average composition. (b) The fraction of complex eigenvalues are plotted as a function of $\alpha$ shows that the fraction increases with $\alpha$ although it does not approach unity.
Figure 6: Pattern formation with decreasing $\Theta$. Panel (a) illustrates the dynamics in the steady state for a fixed randomly sampled realisation of $\mathbb{M}$ and $\mathbb{C}$ for a system with $N=100$ with diagonal chosen from a discrete bimodal distribution. The values chosen for $\Theta$ are indicated in the figure. We generically find spatiotemporal chaos in ten staedy state. (b) illustrates the dynamics in the steady state with the diagonal disorder from an exponential distribution. Each kymograph shows the evolution of the scalar field for one species among $N$ of them.
V Imaginary eigenvalues and steady state dynamics

### v3

@@ III Spinodal
$\Delta$ measures the mean squared deviation from the homogeneous state and vanishes when it is stable. $\Delta$ deviates from zero when the system develops patterns, as seen in the space-time kymograph showing the evolution of the system in one dimension in Fig. 1 (c). Thus, $\Delta$ serves as the order parameter for the transition to the ordered state. The value of $\Delta$ is determined by a spatial and temporal averaging of the numerical results averaging, followed by an ensemble averaging over 5050 realizations of the matrix $\mathbb{M}$, for $N=100$ components. The agreement between theoretical and numerical results in Fig. 2 for both bimodal and exponential distributions (panels (a) and (b)) proves that the stability condition in Eq. (9) which holds for infinitely many components is applicable for a finite number of components as well.
Figure 3: Boundary enclosing the eigenvalues in the complex plane: (a) The boundary and the spectrum of a large random matrix with diagonal disorder drawn from the bimodal distribution, see Eq. (12) for $d_{2}=-d_{1}=d$. (b) The diagonal elements of $\mathbb{C}$ are drawn from a uniform distribution where $d_{-}=1$ and $d_{+}=2$. (c) The diagonal disorder is drawn from an exponential distribution with $d=0.6$. (d) The spectrum and the boundary of support for a large random matrix with diagonal elements from a beta distribution with $a=2$ and $b=5$. For all examples, the random matrix is of the size $\text{N}=1000$ and $\chi=1,\alpha=0.5$. (e) The diagonal disorder are drawn from a trimodal distribution with $d_{1}=0.5$, $d_{2}=2$, $d_{3}=5$, $p_{1}={1}/{2}$ and $p_{2}={1}/{5}$ the random matrix is of a size $N=2000$ and $\chi=1$.
Figure 4: The slope $m$ and the intercept $l_{0}$ for different types of compositional disorder: The slope $m$, referred to in Eq. (9) and (25) for the definition, is calculated for the distributions indicated in the figure. The heatmap of $m$ for the two parameters in (a) and (d) and its variation with the width $d$ in (b) and (c) shows that it is always less than zero. Both the uniform and exponential distributions approach the delta function for $d\to 0$ so that $m$ approaches 1-1 in panels (b) and (c). Panels (e) and (f) show that $l_{0}$ can be positive or negative, meaning that $\Theta_{\rm sp}$ for the passive system can be higher or lower for the corresponding system with identical mean densities.
IV Eigenspectrum and the most unstable eigenvalue
The objective of this section is to establish the results in the previous section with analytical arguments. We first explore the eigenvalue spectrum due to compositions drawn from four different distributions that are qualitatively different from one another, being either bounded or unbounded, and which assume either discrete or continuous values. The boundary enclosing the spectrum of complex eigenvalues is approximated by a continuous curve when $N\to\infty$, as illustrated in Fig. 3 with some examples. In Fig. 3, the markers denote eigenvalues of a single realization of the matrix $\mathbb{D}$ with $N=2000$, while the boundary has been calculated using methods described in [76] which provides an estimate for an infinite number of components. We will briefly describe the analytical approach and use it to justify the form of Eq. (9). The spectral distribution $\rho$ of the eigenvalues $\lambda_{j}$ of $\mathbb{D}$ can be written formally as
@@ IV Eigenspectrum and the most unstable eigenvalue
The boundary of the eigenspectrum of $\mathbb{D}$ (one that encompasses all eigenvalues $\lambda_{j}$ in the complex plane) or equivalently, the support set of $\rho(z)$ is called $\mathcal{S}$ and is given by the points $z\equiv(x,y)$ in the complex plane as indicated in Fig. 3.
@@
As seen in Fig. 3 (a), the bimodal effectively splits the components into two groups containing pNpN and (1p)N(1-p)N species each. For any $p$ and for $d_{1,2}\to 0$ the boundary of the spectrum is simply stretched along the real axis. As $d_{1,2}$ increases, a fissure appears in the boundary, whose location reflects the chosen $p$, being located around $\mbox{Re}(\lambda)=0$ for $p=1/2$. Eventually, at a large enough value of $d$, $\mathcal{S}$ splits into two dis-connected contours, each enclosing pNpN and (1p)N(1-p)N eigenvalues. The numerical results match well with the contour determined from solving Eqs. (18) and (17), as seen in Fig. 3 (a). The distribution can be generalised for multiple discrete values $d_{i}$ with probability $p_{i}$ as follows
$P(\bar{\phi})=\sum_{i=1}^{n}p_{i}\delta(\bar{\phi}-d_{i}),\,\,p_{n}=1-\sum_{i=1}^{n-1}p_{i}.$ (13)
An example of the trimodal distribution and a splitting of the spectrum into three parts in shown in Fig. 3 (e).
The exponential distribution is unbounded and described by a single parameter $d$ that determines the upper-limit of the values that are sampled with high probability.
@@
The boundary in Fig. 3 (b) reflects the asymmetry of the distribution and shows a tail that reflects the large deviations, as only a few species are included with high mean composition.
@@
where $N_{a,b}$ is the normalization constant and $B$ is the Beta function. The method used to calculate the boundary $\mathcal{S}$ is described next.
IV.1 Calculation of the boundary $\mathcal{S}$
The starting point of our analysis are the coupled equations, see Eqs. (17) and (18) below, that determine the contour of points $(x,y)$ in the complex plane enclosing the eigenspectrum $\{\lambda_{j}\}$ of $\mathbb{D}$. To do so, it is essential to introduce the function $\mbox{Re}(g)$ which is related to the resolvent $\sum_{j}(z-\lambda_{j})^{-1}$ (see Appendix VII for details and [76] for the original derivation). The condition that $\mbox{Re}(g)$ assumes finite values outside the boundary $\mathcal{S}$ (i.e. in the region where no eigenvalues appear) and diverges precisely on the boundary was used to determine the relation between the points $(x,y)$ constituting $\mathcal{S}$ and $\mbox{Re}(g)$ [76]. The necessity of introducing $\mbox{Re}(g)$ has been explained in the Appendix VII.
$\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{\sigma^{2}}{\left((\bar{\phi}-x)+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}\right)^{2}+{y^{2}}/{(2\alpha^{2})^{2}}}$
$=1.$ (17)
and
$\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{\bar{\phi}-x+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}}{\left((\bar{\phi}-x)+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}\right)^{2}+{y^{2}}/{(2\alpha^{2})^{2}}}$
$=-\operatorname{Re}\left(g\right),$ (18)
where the integrals are defined over the domain of $\bar{\phi}$. For $P(\bar{\phi})=\delta(\bar{\phi}-\phi_{0})$, the eigenvalues are distributed uniformly with an ellipse with major and minor axes equal to $1-\alpha^{2}$ and $\alpha^{2}$ respectively, such that $\mbox{Re}[\lambda_{0}]=1-\alpha^{2}$ as verified readily from Eqs. (18) and (17), see Appendix VII.1.
The general relation in Eq. (9) is motivated by a close examination of the structure of the coupled equations Eqs. (18) and (17) and verified by numerical calculations using the distributions introduced above. To calculate $\mathcal{S}$ for any distribution $P$, we define scaled and shifted coordinates $\tilde{x}$ and $\tilde{y}$
@@ IV.1 Calculation of the boundary $\mathcal{S}$
Notice that Eqs. (17) and (18) rewritten using $\tilde{x}$ and $\tilde{y}$ are independent of $\alpha$, a property that will be used in evaluating $\mathcal{S}$. As a first step to calculating $\mathcal{S}$ we use Eq. (19) to re-write Eq. (17) as
[…]
thus defining the function $h_{\tilde{y}}$. The function $h_{\tilde{y}}$ can in principle be inverted to obtain $\tilde{x}$, for a fixed $\tilde{y}$ as
[…]
In the second step, $\mbox{Re}(g)$ is calculated using Eq. (18),
$\mbox{Re}(g)[\tilde{x},\tilde{y}]=-\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{\bar{\phi}-\tilde{x}}{(\bar{\phi}-\tilde{x})^{2}+{\tilde{y}^{2}}}.$ (22)
The following expression for $x$ holds for a given $y$
[…]
$+(1-2\alpha^{2})\sigma^{2}\mbox{Re}(g)[h^{-1}(\sigma^{-2};{\tilde{y}}),\tilde{y}$], (23)
showing that $x$ depends explicitly on $\alpha$ only through the second term in the R.H.S. of Eq. (IV.1). The procedure yields the set of points $(x,y)$ defining the curve $\mathcal{S}$ that serves well as the boundary for the eigenvalues plotted for $N=2000$ in the complex plane, see Fig. 3.
$\alpha=1/\sqrt{2}$ is a special value of $\alpha$ because at this point $\tilde{x}=x$, and it is particularly easy to carry out the algebraic operations described above, a point illustrated with the example of the bimodal distribution in Appendix VII. The physical meaning of $\tilde{x}$ and $\tilde{y}$ is now clear - they are simply the curve $\mathcal{S}$ at $\alpha=1/\sqrt{2}$. Consider now two values of $\alpha$, $\alpha_{1,2}$ both different from $1/\sqrt{2}$, at fixed $y$ we have using Eq. (19)
$\frac{x(\alpha_{1})-x(\alpha_{2})}{\alpha_{2}^{2}-\alpha_{1}^{2}}=2\left[\mbox{Re}(g)(\tilde{x},\tilde{y})\right]\sigma^{2}$
$\frac{y(\alpha_{1})}{y(\alpha_{2})}=\frac{\alpha_{2}^{2}}{\alpha_{1}^{2}}.$ (24)
Eq. (IV.1) suggests that the curve $\mathcal{S}$ can be calculated at all values of $\alpha$ if we calculate it at $\alpha=1/\sqrt{2}$.
@@
Notice now that for two different values of $\alpha$, the $y$ values of the curves $\mathcal{S}$ are proportional. For $\tilde{y}=0$, i.e., the points at which $\mathcal{S}$ intersects the $x$ axis are the same for $\mathcal{S}$ for any value of $\alpha$. This implies that the topology of $\mathcal{S}$, meaning the number of times the curve $\mathcal{S}$ intersects the real axis, is determined by $\{D\}$ only and does not depend on $\alpha$ at all, as illustrated using the bimodal distribution in panel (e) of Fig. 3. The curve in $\mathcal{S}$ splits into two above a threshold value of $d$, as seen in Fig. 3 (a), meaning that $\tilde{y}$ vanishes at four points on the $x$-axis. The topology of the boundary is preserved for all values of $\alpha$ as shown in Fig. 3 (e) for a trimodal distribution where the $\{D\}$ are fixed and $\alpha$ only is varied. The spectrum for the values indicated at the values of alpha indeed follows the insight from theory.
IV.2 Expression for $\lambda_{0}$
By the definition of $\lambda_{0}$ (intercept $x$ of the curve $\mathcal{S}$ at $y=0$), we have using Eq. (IV.1)
$\lambda_{0}(\alpha)=L_{0}-m\alpha^{2},$ (25)
where the intercept $L_{0}$ that determines the threshold of instability for the passive system is given by
$L_{0}=h^{-1}\left(\sigma^{-2};0\right)+\sigma^{2}\mbox{Re}(g)[h^{-1}\left(\sigma^{-2};0\right),0].$ (26)
$\lambda_{0}$ varies quadratically with $\alpha^{2}$ with the constant of proportionality or ’slope’ $m$ given by
$m=2\sigma^{2}\mbox{Re}(g)[h_{0}^{-1}\left(-\sigma^{-2}\right),0].$ (27)
Substituting Eq. (27) in the definition of the threshold of instability $\Theta_{\rm{sp}}=-1-\mbox{Re}[\lambda_{0}]$ produces the expression in Eq. (9) with the identification $l_{0}=-1-L_{0}$.
@@ IV.2 Expression for $\lambda_{0}$
Figure 5: Statistics of $\lambda_{0}.$ In the limiting case of infinite conserved densities, $\lambda_{0}$ is real. For any finite-sized system, $\langle\lambda_{0}\rangle$ averaged over ensembles of $\mathbb{D}$, is in fact complex. (a) The imaginary and real parts, $\mbox{Im}(\lambda_{0})$ and $\mbox{Re}(\lambda_{0})$, are plotted as a function of $\alpha$, encoded in the colour of the marker. The error bar indicates the variance. At each value of $\alpha$, we consider $10^{4}$ realizations of $\mathbb{D}$ for $N=300$. From the plots, it is clear that the variability in $\mbox{Im}(\lambda_{0})$ depends on the type of compositional disorder. Notice also that the variations are increased compared to the uniform average composition. (b) The fraction of complex eigenvalues is plotted as a function of $\alpha$, showing that the fraction increases with $\alpha$ although it does not approach unity.
From Eq. (27), it is evident that $m$ depends only on the parameters $\{D\}$ meaning that it is enough to resolve the sign of $m$ for any non-zero value of $\alpha$, which following the train of discussion so far in the section is accomplished easily at $\alpha=1/\sqrt{2}$.
The form of the diagonal terms that we consider represents squared deviations of the mean density of each species from the average and thus they are always positive quantities bounded by zero from below. From Eq. (19) it is clear that at the extreme value $\lambda_{0}$ of $\tilde{x}=x$,
$m=-2\sigma^{2}\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{1}{(\bar{\phi}-\lambda_{0}(1/\sqrt{2}))}$ (28)
the sign of the integrand in Eq. (22) is determined by the sign of $(\bar{\phi}-\lambda_{0}(1/\sqrt{2}))$, a quantity whose sign cannot change in the domain of $\bar{\phi}$ for a distribution with a well-defined $\lambda_{0}$. This argument constraints the signs of the deviation $\bar{\phi}-\lambda_{0}({1/\sqrt{2}})$ and thus $m$ to be either always negative and positive respectively, or vice versa. The two cases represent the two extreme bounds of the eigenvalues, and clearly $m>0$ corresponds to the eigenvalue with the smallest real part.
To test the validity of the theoretical predictions, we plot $m$ by varying the parameters $\{D\}$ for the distribution shown in Fig. 1. For each value indicated in the axes of the panels in Fig. 4, we calculate $\lambda_{0}$ at two values of $\alpha^{2}$ and calculate the slope and verify that it is indeed positive everywhere (see panels (a-d) of Fig. 4).
$m=\frac{\lambda_{0}(\alpha_{1})-\lambda_{0}(\alpha_{2})}{\alpha^{2}_{2}-\alpha^{2}_{1}}.$ (29)
We also calculate $L_{0}$ and find that it can assume both positive and negative signs depending on the parameters specifying the distribution meaning that the stability of the passive system, ($\alpha=0$) does not follow a universal trend. Panels (e) and (f) of Fig. 4 shows the variation in $\Theta_{\rm{sp}}(0)=l_{0}=-L_{0}-1$, (refer to Eq. (9)) for the bimodal and exponential distributions respectively.
Figure 6: Steady state dynamics for mixture (composition drawn from a bimodal distribution): The left panels illustrate the dynamics in the steady state for randomly sampled realizations of $\mathbb{M}$ and $\mathbb{C}$ for a mixture with $N=100$ components. The average density of each species is chosen from a discrete bimodal distribution. i.e. it can attain two values either $d_{1}$ or $d_{2}$. The effective temperature $\Theta$ is indicated in the figure. For $\Theta$ very close to $\Theta_{\rm{sp}}$, the system undergoes bulk phase separation. As seen in the histograms of the densities of a few species selected at random, shows that species 242-4 exhibit more than two peaks. This means that the system has compartmentalised into more than two compartments, a feature common in passive systems. At larger $\Theta$, there are two asymmetric peaks that get smoother with decreasing $\Theta$.
V Imaginary eigenvalues and steady state patterns